# SURT 2.0: Advances in Transducer-based Multi-talker Speech Recognition

### v1

@@ VI-A Comparison with baselines
First, we compared the SURT 2.0 base and large models with the baselines from literature: SURT [26] and MT-RNNT [27], on the LibriCSS “anechoic” and “replayed” recordings. These comparisons are shown in Table III. The models were trained on LSMix-full, and adapted to LibriCSS using the dev set. For adaptation, we segmented the dev sessions in 2 ways, by cutting at maximum pause durations of 0.1s and 0.5s, respectively. This resulted in 381 sub-sessions of average duration 17.9s, totalling approximately 1.88h each from anechoic and replayed conditions and containing 18.6% overlapped speech. The model was then trained on the combined adaptation data for 8 epochs (for base) or 15 epochs (for large). The learning rate was warmed up to 0.0004 for 2 (or 4) epochs and decayed thereafter. A single V100 GPU was used for adaptation.
@@ VI-C Effect of auxiliary objectives
First, we see that adding CTC loss on the encoder improved WER mainly for highly overlapped sessions. We obtained relative improvements of 10.2% and 11.1% on the OV30 and OV40 sessions, respectively. For sessions with more overlaps and turn-taking, both output branches may contain several speech segments. We conjecture that an auxiliary CTC objective may be useful in aligning the segments to the corresponding audio during training, resulting in better modeling for high overlap sessions. To validate our conjecture, we plotted the the occupation probabilities for the nodes in the RNN-T lattice (of shape $T\times U$), as obtained from the gradients of the simple additive joiner used in the pruned transducer loss [88]. These values should correspond to a soft alignment between the input and the label sequence [^7] It makes more sense to use this value instead of CTC alignments because (i) the model trained without $\mathcal{L}_{\text{ctc}}$ cannot provide corresponding alignments, and (ii) we use the transducer head for the ASR task.. In Fig. 4, we show example plots for a randomly selected mixtures from the training set, using models trained with and without $\mathcal{L}_{\text{ctc}}$. When the auxiliary CTC loss was used, the model was able to better align the silence region (time frames 100 to 250) in channel 2, as discernible through the bright horizontal line.
@@ VI-E Measuring leakage and omission
If emission time-stamps are known, these definitions may be modified to include time windows for n-gram search. In their absence, the leakage@n and omission@n rates are upper and lower bounds on the actual leakage and omission, respectively. Nevertheless, by quantifying these error types explicitly, we can gain some insights about the model behavior.
@@ Acknowledgments
We thank Aparna Khare for discussions and insights on multi-turn RNN-T, and Piotr Zelasko, Fangjun Kuang, and Thilo von Neumann for help with Lhotse, Icefall, and meeteval, respectively.

### v2

@@ VI-A Comparison with baselines
First, we compared the SURT 2.0 base and large models with the baselines from literature: SURT [27] and MT-RNNT [28], on the LibriCSS “anechoic” and “replayed” recordings. These comparisons are shown in Table III. We cannot compare SURT with published results on t-SOT [24] since the latter is evaluated using the asclite-based speaker-agnostic WER (SAg-WER) metric. SAg-WER requires token-level time-stamp estimation, and does not penalize an utterance being split into multiple “channels,” which is penalized by ORC-WER. The models were trained on LSMix-full, and adapted to LibriCSS using the dev set. For adaptation, we segmented the dev sessions in 2 ways, by cutting at maximum pause durations of 0.1s and 0.5s, respectively. This resulted in 381 sub-sessions of average duration 17.9s, totalling approximately 1.88h each from anechoic and replayed conditions and containing 18.6% overlapped speech. The model was then trained on the combined adaptation data for 8 epochs (for base) or 15 epochs (for large). The learning rate was warmed up to 0.0004 for 2 (or 4) epochs and decayed thereafter. A single V100 GPU was used for adaptation.
@@ VI-C Effect of auxiliary objectives
First, we see that adding CTC loss on the encoder improved WER mainly for highly overlapped sessions. We obtained relative improvements of 10.2% and 11.1% on the OV30 and OV40 sessions, respectively. For sessions with more overlaps and turn-taking, both output branches may contain several speech segments. We conjecture that an auxiliary CTC objective may be useful in aligning the segments to the corresponding audio during training, resulting in better modeling for high overlap sessions. To validate our conjecture, we plotted the occupation probabilities for the nodes in the RNN-T lattice (of shape $T\times U$), as obtained from the gradients of the simple additive joiner used in the pruned transducer loss [89]. These values should correspond to a soft alignment between the input and the label sequence [^7] It makes more sense to use this value instead of CTC alignments because (i) the model trained without $\mathcal{L}_{\text{ctc}}$ cannot provide corresponding alignments, and (ii) we use the transducer head for the ASR task.. In Fig. 5, we show example plots for a randomly selected mixtures from the training set, using models trained with and without $\mathcal{L}_{\text{ctc}}$. When the auxiliary CTC loss was used, the model was able to better align the silence region (time frames 100 to 250) in channel 2, as discernible through the bright horizontal line.
@@ VI-E Measuring leakage and omission
If emission time-stamps are known, these definitions may be modified to include time windows for n-gram search. In their absence, the leakage@n and omission@n rates are upper and lower bounds on the actual leakage and omission, respectively. Nevertheless, by quantifying these error types explicitly, we can gain some insights about the model behavior. Our SURT model in this paper does not predict token time-stamps, so we estimate omission and leakage over the entire hypotheses.
@@ Acknowledgments
We thank Aparna Khare for discussions and insights on multi-turn RNN-T, and Piotr Zelasko, Fangjun Kuang, and Thilo von Neumann for help with Lhotse, Icefall, and meeteval, respectively. We are grateful to the anonymous reviewers for their feedback.